Ask
your own question, for FREE!
Mathematics
11 Online
You have a 3-card deck containing a king, a queen, and a jack. You draw a random card, then without putting it back you draw a random second card from the ones that are left. Use a tree diagram to calculate the probability that you draw exactly 1 jack.
Still Need Help?
Join the QuestionCove community and study together with friends!
First probability: 2/3 (you can take either a K or Q) Second probability: 1/2 (you can take either a K or Q, depending on which one is left) 2/3*1/2 = 1/3
@daniel I think you got it wrong. ..... What abt the case when jack can be drawn first, and in the second draw there can be a king or queen? I think the correct answer would be (1/3)*1 + (2/3)*(1/2) = (1/3)+(1/3)=2/3
Woops: I really messed up on this one: I read it so that you had to be left with one Jack. You are 100% correct: 2/3
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
9 hours ago
0 Replies
0 Medals
4 days ago
3 Replies
2 Medals
1 week ago
3 Replies
0 Medals
1 day ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal